F(x)= (3x+2)/(x+7)
P=(1, 5/8)
Equation of the tangent line to the graph of f at p.

Answer by Boreal(15235)

f(x)=(3x+2)/(x+7)
derivative is 19/(x+7)^2, and when x=1, derivative is 19/64
point slope formula for line where y-y1=m(x-x1), (x1, y1) is point
y-(5/8)=(19/64)(x-1)=(19/64)x-(19/64)
y=(19/64)x+(21/64)